Problem 12-08 Bonds of anca Corporation will per value of $1,000 000 mature in four years, and current yield

Bonds of Francesca Corporation with a par value of $1000 sell for $980, mature in four years, and have a 8% annual coupon rate period semianually.

a) current yield ______ %annually
b) yield to maturity, to the nearest basis point ______ %annually
c) horizon yield (or realized return) for an investor with a two-year holding period and a reinvestment rate of 6% over the period. At the end 2 years, the 8% coupon bonds with 2 years remaining will sell to yield 8%. _____% annually
Answer #1

a) Current Yield = Coupon / Price = 8% x 1000 / 980 = 8.16%

b) YTM can be calculated using I/Y function

N = 4 x 2 = 8, PMT = 40, PV = -980, FV = 1000

=> Compute I/Y = 4.30% (semi-annual)

Annual YTM = 4.30% x 2 = 8.60%

c) After two years, the price of bond = $1000 as coupon rate is equal to YTM.

Future Value of reinvestment of coupon can be calculated using FV.

N = 2 x 2 = 4, PMT = 40, PV = 0, I/Y = 6%/2 = 3% => Compute FV = $167.35

Realized return can be calculated using I/Y function

N = 2, PV = -980, FV = 167.35 + 1000 = 1167.35, PMT = 0

=> Compute I/Y = 9.14%